In a population of 4,549 moths, you observe 622 moths that have large spots on their wings. You identify that having large spots is dominant over small spots (aa), but the trait displays incomplete dominance (Aa individuals have medium spots). What is the frequency of the A allele in the population (be sure to round to 4 decimal places)?
Answer #1

Out of 4549, 622 moths have large spots on their wings

3927 moths have both small and medium spots

Allele frequencies p+q=1

p=no. of copies of A allele/total population number

p=1244/4549 = 0.2734

q=1-p => 1-0.2734 = 0.7266